OpenStudy (anonymous):

Explain the steps you would use to compare four-fifteenths and five-sixths. Remember, you can use the forward-slash key (/) to type fractions. For example, type one-fourth as one slash four.

OpenStudy (jdoe0001):

so, ahemm. what do we do again? compare them... how ?

OpenStudy (anonymous):

what method would you use @jdoe0001

OpenStudy (jdoe0001):

well, I'd make them both the same denominator, that is the same whole how? well, multiply the both denominator and numerator by the denominator of the other thus  \(\bf \cfrac{4 \times 6}{15 \times 6}, \cfrac{5 \times 15}{6 \times 15} \implies \cfrac{24}{90}, \cfrac{75}{90}\)